### v1

@@ A.1 Implementation details
Table 3: Training hyperparameters
A.2 Language specific knowledge
Beyond measuring similarities between languages using the Jaccard Index, we also analyze differences by identifying answers that are known in language A but unknown in language B. This allows us to examine the distribution of languages within the 40% of answers that are not correctly predicted by both languages. The results, presented in Figure 5, reveal an almost symmetrical distribution of known and unknown answers across most language pairs. However, notable deviations occur for languages with significantly lower overall accuracy. Specifically, Hindi shows a greater disparity in the Llama-3.2 models, while both Hindi and Portuguese exhibit this trend in the Llama-3.1-8B model.
@@ A.3 Parameter update
A.4 Partial Training
Due to space limitations, the results of partial training on Italian, Portuguese, and Spanish are provided in Figure 7. Additionally, the CoCo-CoLa ratios for both partially trained and fully trained models are shown in Table 4 for Llama-3.2-1B, Table 5 for Llama-3.2-3B, and Table 6 for Llama-3.1-8B. These comparisons highlight the consistently superior CoCo-CoLa ratio in the partial training of final layers.

### v2

@@ A.3 Implementation details
A.4 Language specific knowledge
Beyond measuring similarities between languages using the Jaccard Index, we also analyze differences by identifying answers that are known in language A but unknown in language B. This allows us to examine the distribution of languages within the 40% of answers that are not correctly predicted by both languages. The results, presented in Figure 5, reveal an almost symmetrical distribution of known and unknown answers across most language pairs. However, notable deviations occur for languages with significantly lower overall accuracy. Specifically, Hindi shows a greater disparity in the Llama-3.2 models, while both Hindi and Portuguese exhibit this trend in the Llama-8B model.
@@ A.5 Parameter update
A.6 Partial Training
Due to space limitations, the results of partial training on Italian, Portuguese, and Spanish are provided in Figure 7. Additionally, the CoCo-CoLa ratios for both partially trained and fully trained models are shown in Table 5 for Llama-1B, Table 6 for Llama-3B, and Table 7 for Llama-8B. These comparisons highlight the consistently superior CoCo-CoLa ratio in the partial training of final layers.
@@ A.6 Partial Training
Table 8: CoCo-CoLa Ratios (%) for different languages across finetuned Gemma-3-4B models.
[…]
French 19.22 64.26 89.99 90.14
German 15.23 86.70 93.03 92.64
Hindi 9.39 92.74 96.30 97.19
Italian 14.84 85.03 91.20 91.08
Portuguese 14.98 70.93 88.40 90.69
Spanish 14.70 68.14 90.19 91.36